# Efficient Analysis of Latent Spaces in Heterogeneous Networks

### v2

@@ 3 Estimation of the latent spaces
In the following, we assume observed data follow the model (3) with parameters $(Z,W_{t})=(Z^{\star},W_{t}^{\star})$ over $1\leqslant t\leqslant T$. We aim to estimate the underlying true parameters $Z^{\star}$ and $W^{\star}_{t}$ from the observed data. To address all the challenges discussed in Section 2.3, we propose a three-stage estimation procedure:
@@
As Step (a) analyzes each network individually, $Y_{t}^{\star}$ can only be identified up to an orthogonal group transformation $\mathcal{O}(d_{t})$ under (3). Consequently, column spaces of $Z^{\star}$ and $W_{t}^{\star}$ cannot be separated, necessitating the joint analysis across $T$ networks in Steps (b) and (c). As Section 2.2 suggests, $Z^{\star}$ and $W_{t}^{\star}$ can only be estimated up to factor-specific orthogonal group transformations $\mathcal{O}(k)$ and $\mathcal{O}(k_{t})$, respectively. For simplicity, we refer to estimating $Z^{\star}$ and $W_{t}^{\star}$ without explicitly emphasizing the factor-specific transformations when there is no ambiguity. In the following, we assume the latent dimensions $k$ and $k_{t}$’s are given, and discuss how to consistently estimate them in Section B of the Supplementary Material.
@@ 3.1 Individual estimation
We first estimate each $Y_{t}^{\star}$ up to an orthogonal group transformation, which can be achieved by examining each single network individually. For a single network, there have been a variety of studies to estimate latent spaces for both random dot product models and latent space models with non-linear link function (Chatterjee,, 2015; Ma et al.,, 2020; Zhang et al.,, 2020). For the subsequent analysis in Sections 3.2 and 3.3, it suffices to obtain an estimator $\mathring{Y}_{t}$ satisfying $\mathrm{dist}^{2}(\mathring{Y}_{t},Y_{t}^{\star})$ being of the order of $O_{p}(1)$ up to logarithmic factors.
The desired error rate is consistent with existing results in the literature and may be achieved by various methods. In this work, we adopt a strategy that maximizes the likelihood function of $Y_{t}^{\star}$ via projected gradient descent for its simplicity and computational efficiency. The method generalizes that in Ma et al., (2020) for a binary network to models under a general exponential family distribution. As we consider a general link function in the model (3), different adjustments for links are needed are in both methodology and theory. We defer the details to Section A in the Supplementary Material. Notably, any other methods that can estimate $Y_{t}^{\star}$ with the desired error rate can also be used in our subsequent analysis.
@@ 3.2 Shared space hunting
Although estimating $Y_{t}^{\star}=[Z^{\star},W_{t}^{\star}]$ from each network is well-established, $Y_{t}^{\star}$ is only identifiable up to an unknown orthogonal transformation under a single-network model. As a result, $Z^{\star}$ cannot be directly obtained. To address this issue, we leverage joint information of $Y_{t}^{\star}$’s and develop a novel and efficient spectral method below. To motivate our proposed construction, we first introduce an algebraic result in an oracle scenario below, where accurate estimates $Y_{t}$’s that equal $Y_{t}^{\star}$’s up to unknown orthogonal transformations are available.
@@
    there exist $1\leqslant t<s\leqslant T$ such that the columns of $[Z^{\star},W_{t}^{\star},W_{s}^{\star}]$ are linearly independent.
Let $\mathcal{T}=\{(t,s):1\leqslant t<s\leqslant T\text{ and assumption (ii) holds under }(t,s)\}$. Then for $(t,s)\in\mathcal{T}$,
@@
Assumption (i) in Proposition 2 posits an oracle situation where we can accurately estimate each $Y_{t}^{\star}$ individually up to the orthogonal group transformation $\mathcal{O}(d_{t})$. Assumption (ii) in Proposition 2 is the same as the assumption (ii) in Proposition 1 for the model identifiability. Eq. (4) provides the theoretical foundation for identifying $Z^{\star}$ based on joint information from $Y_{t}$ and $Y_{s}$, which is unachievable when only investigating each $Y_{t}$ individually. However, (4) may not be directly applicable to estimating $Z^{\star}$ in practice due to the following reasons. First, an estimator $\mathring{Y}_{t}$ obtained in Step 1 may not exactly satisfy $\mathring{Y}_{t}\mathring{Y}_{t}^{\top}=Y_{t}^{\star}Y_{t}^{\star\top}$, and consequently, (4) may not hold precisely. Second, (4) relies on the index set $\mathcal{T}$, whose definition depends on the unknown true parameters.
To address these issues, we develop practical procedures based on spectral information of the estimates $\mathring{Y}_{t}$’s in Step 1, which can be shown to approximate the true parameters $Y_{t}^{\star}$’s well under suitable conditions. To estimate $Z^{\star}$, we first construct a screening method to estimate $\mathcal{T}$. Let $\sigma_{l}(B)$ denote the $l$-th largest singular value af a matrix $B$, and define $\mathcal{R}_{l,j}(B)=\sigma_{l}(B)/\sigma_{j}(B)$. Then we have
@@ 3.3 Likelihood-based refinement
$\ell(Z,W)\propto\sum_{t=1}^{T}\sum_{1\leqslant i\leqslant j\leqslant n}\left\{A_{t,ij}\big{(}\langle z_{i},z_{j}\rangle+\langle w_{t,i},w_{t,j}\rangle\big{)}-\nu\big{(}\langle z_{i},z_{j}\rangle+\langle w_{t,i},w_{t,j}\rangle\big{)}\right\},$ (6)
where we denote $W=[W_{1},\ldots,W_{T}]$ for simplicity. Our refinement procedure first conducts the projected gradient descent (Chen and Wainwright,, 2015), primarily using first-order derivatives of the likelihood function that can be efficiently computed. Specifically, the algorithm descends parameters along their gradient directions with pre-specified step sizes and then projects updated estimates to pre-specified constraint sets for parameters. The detailed steps are summarized in Algorithm 2 with notations explained in Section 3.4. Let $\check{Z}$ and $\check{W}_{t}$’s denote the estimates after sufficient iterations of the projected gradient descent.
Then the refinement procedure further constructs a second-order update, which utilizes efficient influence function and provides convenience for theoretically establishing desired statistical error rate. For the ease of introducing formula below, we denote a vectorization of all the entrywise parameters of $[Z,W_{1},\ldots,W_{T}]$ as
@@
where $k_{\operatorname{sum}}=k_{1}+\cdots+k_{T}$, and similarly define $\check{v}$ as the vectorization of $[\check{Z},\check{W}_{1},\ldots,\check{W}_{T}]$. Then $\ell(Z,W)=\ell(v)$, and let ˙$\dot{\ell}(v)$ denote the partial derivative of $\ell(v)$ with respect to $v$. We construct a second-order update as
@@
Our proof shows $I(\check{v})$ in (8) is singular, necessitating the use of pseudo inverse in (8). Intuitively, this singularity arises from the unidentifiability of latent factors $Z$ and $W_{t}$ (Little et al.,, 2010). It makes the theoretical analysis challenging and conclusions for classical one-step estimator (Van der Vaart,, 2000) inapplicable. Moreover, the efficient one-step update (8) is simultaneously constructed for $1+T$ groups of latent factors $[Z,W_{1},\ldots,W_{T}]$. This markedly differs from He et al., (2023) that constructs one-step estimator for shared latent factors only and requires network-specific parameters to be identifiable. As a result, our theoretical analysis pioneers a new approach that jointly analyzes all the parameters in (7) and tackles inherent singularity of the information matrix from multiple groups of latent factors. More details are given in Section D.5.4 of the Supplementary Material.
@@ 3.4 Algorithm details
10Construct $\hat{Z}$ and $\hat{W}$ as the matrix version of (8) with $(\check{Z},\check{W})=(Z^{R},W^{R})$.
@@
In Algorithm 1: The threshold $\tau_{1}$ is chosen to select index pairs $(t,s)\in\mathcal{T}$. As Section 3.2 suggests, the choice of $\tau_{1}$ depends on statistical properties of the input initial estimates $\mathring{Y}_{t}$ in general. In this paper, we construct $\mathring{Y}_{t}$ by the method in Section 3.1, whose estimation errors will be established in Section 4. To achieve theoretical guarantee, it suffices to set the threshold to satisfy $1\ll\tau_{1}\lesssim\sqrt{\log n}$. In our numerical implementation, we set $\tau_{1}=\sqrt{2\log n}$.
In Algorithm 2: During the gradient descent, we let ˙$\dot{\ell}_{Z}(Z,W)\in\mathbb{R}^{n\times k}$ and ˙$\dot{\ell}_{W_{t}}(Z,W)\in\mathbb{R}^{n\times k_{t}}$ represent the partial derivatives of $\ell(Z,W)$ with respect to matrices $Z$ and $W_{t}$, respectively. Let $\mathcal{P}_{\mathcal{C}}(\cdot)$ denote a projection operator given a pre-specified constraint set $\mathcal{C}$, and we consider two constraint sets
@@
for $Z$ and $W_{t}$, respectively. Eq. (9) constraints the two-to-infinity norms of $Z$ and $W_{t}$ to be bounded by $M_{1}$, which also corresponds to the constraints on the true parameters $(Z^{\star},W_{t}^{\star})$ to be specified Condition 2 in Section 4 below. It is worth mentioning that the projection step in Algorithm 2 is required for the convenience of the proof, and it is skipped in practical implementation, which can give equally good numerical performance. To achieve theoretical guarantee, we choose step sizes $\eta_{Z}=\eta/(nT)$ and $\eta_{W}=\eta/n$ for some small constant $\eta>0$, and the number of iterations $R\gg\log(nT)$. In our numerical implementation, we use Barzilai-Borwein step sizes (Barzilai and Borwein,, 1988) and set $R=1000$.
Condition 1 summarizes the requirements on the hperparameters in Algorithms 1 and 2 to establish theoretical results in Section 4.
@@
Assume the hyperparameters in Algorithms 1 and 2 satisfy: (i) $1\ll\tau_{1}\lesssim\sqrt{\log n}$. (ii) Step sizes $\eta_{Z}=\eta/(nT)$ and $\eta_{W}=\eta/n$ for some small constant $\eta>0$. (iii) Constraint sets $\mathcal{C}_{Z}$ and $\mathcal{C}_{W_{t}}$ are chosen as in (9). (iv) The number of iterations $R\gg\log(nT)$.
@@ 3.5 Related works
The proposed estimation scheme in Section 3 shares connections with, yet notably differs from, existing methods. The shared space hunting Algorithm 1 is developed based on spectral properties of latent factors. While spectral-based analysis is common in network studies (Rohe et al.,, 2011; Chatterjee,, 2015; Arroyo et al.,, 2021), our procedure utilizes unique properties of multiple networks to separate the shared and distinct latent spaces, and thus cannot be directly implied by existing methods. The likelihood refinement Algorithm 2 utilizes the projected gradient descent, which is known to be computationally efficient for maximizing the non-convex likelihood functions for both single network (Ma et al.,, 2020) and multiple networks (Zhang et al.,, 2020). Nevertheless, the oracle estimation error rates under our model cannot be directly attained through existing technical developments of those methods. To address the theoretical challenges, Algorithm 2 further constructs a novel second-order update estimator (8) for multiple heterogeneous networks. The construction is motivated from the classical one-step estimator, typically obtained through a linear approximation of score equation (Van der Vaart,, 2000). Similar idea has also been recently used to construct efficient estimators for latent factors in network analysis, including the random dot product model for a single network (Xie and Xu,, 2023) and a semiparametric longitudinal network model (He et al.,, 2023). However, our model (1) substantially differs from those models in that it embraces a general edge distribution and also allows distinct latent factors over multiple networks. The extended flexibility posits unique challenges in identifying shared spaces and obtaining efficient estimation errors, necessitating novel developments of efficient score updating formula and theory.
The studied model (1) is motivated from the MultiNeSS model proposed by MacDonald et al., (2022), which similarly considers (1) and assumes that part of the latent vectors are shared across the networks, i.e., $y_{t,i}=(z_{i}^{\top},w_{t,i}^{\top})^{\top}$. The difference is that in their work $\Theta_{t,ij}=\kappa(y_{t,i},y_{t,j})$ with the similarity function $\kappa(x,y)=x^{\top}\mathrm{I}_{p,q}y$, where $\mathrm{I}_{p,q}$ represents the block-diagonal matrix with $\mathrm{I}_{p}$ and $-\mathrm{I}_{q}$ on the upper left and lower right blocks, respectively. Although a more general similarity function is proposed, their theoretical analysis is restricted to Gaussian distribution with identity link, which can be inappropriate for network data with binary edges or count edges. Alternatively, we will next establish theoretical results that encompass various exponential family distributions. Please also see Remark 3 for more discussions on the comparison of technical conditions and results.
@@
Model Link $g(x)$ ER-Shared ER-Individual Constraints on $(n,T)$
Arroyo et al., (2021) $x$ $O_{p}\biggr{(}\frac{1}{n}+\frac{1}{T}\biggr{)}$- --
MacDonald et al., (2022) $x$ $O_{p}\biggr{(}\frac{1}{T}\biggr{)}$ $O_{p}(1)$ $T=o(n^{\frac{1}{2}})$
Zhang et al., (2020) $\operatorname{logit}(x)$ $O_{p}\biggr{(}1+\frac{1}{T}\biggr{)}$- $T=O(n)$
He et al., (2023) $\log(x)$ $O_{p}\biggr{(}\frac{1}{n}+\frac{1}{T}\biggr{)}$ $O_{p}(1)$-
This work General $O_{p}\biggr{(}\frac{1}{n}+\frac{1}{T}\biggr{)}$ $O_{p}(1)$-
(Error rates above are simplified to emphasize their orders with respect to $n$ and $T$ up to logarithmic factors and also rescaled to be comparable. “--” represents inapplicable information. See more details in Section F of the Supplementary Material.)
Besides the studied model (1), there are various other models extracting node-wise latent embeddings from multiple networks. One research line extends the random dot product graphs (Young and Scheinerman,, 2007; Athreya et al.,, 2018) to multilayer networks (Arroyo et al.,, 2021; Zheng and Tang,, 2022). These models typically corresponds to an identity link function in (3) and thus cannot directly address the challenges under models with general non-linear link functions. Another class of models adopt non-linear canonical link functions (Hoff et al.,, 2002; Ma et al.,, 2020) and develop extensions for multiple networks (Zhang et al.,, 2020; MacDonald et al.,, 2022; He et al.,, 2023). Our developments under model (1) advance the literature by establishing a unified theoretical framework that can encompass a variety of general link functions and allow heterogeneous latent factors with complex relationships. We summarize properties and results under the related models in Table 1, where the theory under (1) will be formally presented in Section 4. Notably, oracle error rate of the shared parameters in this paper is consistent with earlier work in Arroyo et al., (2021), MacDonald et al., (2022), and He et al., (2023).
Targeting at distinct properties and challenges, researchers have also proposed other models for multiple networks. Examples include multilayer stochastic block models for community detection (Han et al.,, 2015; Paul and Chen,, 2016; Lei et al.,, 2019), and tensor-decomposition-based models for effective dimension reduction (Lyu et al.,, 2023; Zhang and Wang,, 2023), etc. Although these models may also incorporate shared and heterogeneous structures across multiple networks, their analytical objectives and challenges differ from our focus on estimating continuous node-wise latent embeddings and thus cannot be directly compared as in Table 1.

### v3

@@ 3 Estimation of the latent spaces
In the following, we assume the observed data follow the model (2) with parameters $(Z,W_{t})=(Z^{\star},W_{t}^{\star})$ over $1\leqslant t\leqslant T$. We aim to estimate the underlying true parameters $Z^{\star}$ and $W^{\star}_{t}$ from the observed data. To address all the challenges discussed in Section 2.3, we propose a three-stage estimation procedure:
@@
As Step (a) analyzes each network individually, $Y_{t}^{\star}$ can only be identified up to an orthogonal group transformation $\mathcal{O}(d_{t})$ under (2). Consequently, column spaces of $Z^{\star}$ and $W_{t}^{\star}$ cannot be separated, necessitating a joint analysis across $T$ networks in Steps (b) and (c). As Section 2.2 suggests, $Z^{\star}$ and $W_{t}^{\star}$ can only be estimated up to factor-specific orthogonal group transformations $\mathcal{O}(k)$ and $\mathcal{O}(k_{t})$, respectively. For simplicity, we refer to estimating $Z^{\star}$ and $W_{t}^{\star}$ without explicitly emphasizing these group transformations when there is no ambiguity. In the following, we assume the latent dimensions $k$ and $k_{t}$’s are fixed and given, and discuss how to consistently estimate them in Section B of the Supplementary Material.
@@ 3.1 Individual estimation
We first estimate each $Y_{t}^{\star}$ up to an orthogonal group transformation, which can be achieved by examining each single network individually. For a single network, there have been extensive studies to estimate latent spaces for both random dot product models and latent space models with non-linear link function (Chatterjee,, 2015; Ma et al.,, 2020; Zhang et al.,, 2020). For the subsequent analysis in Sections 3.2 and 3.3, it suffices to obtain an estimator $\mathring{Y}_{t}$ satisfying $\mathrm{dist}^{2}(\mathring{Y}_{t},Y_{t}^{\star})$ being of the order of $O_{p}(1)$ up to logarithmic factors.
The desired error rate is consistent with existing results in the literature and may be achieved by various methods. In this work, we adopt a strategy that maximizes the likelihood function of $Y_{t}^{\star}$ via projected gradient descent for its simplicity and computational efficiency. The method generalizes that in Ma et al., (2020) for a binary network to models under a general edge distribution. As we consider a general link function in the model (2), different adjustments for links are needed are in both methodology and theory. We defer the details to Section A in the Supplementary Material. Notably, any other methods that can estimate $Y_{t}^{\star}$ with the desired error rate can also be used in our subsequent analysis.
@@ 3.2 Shared space hunting
Although estimating $Y_{t}^{\star}=[Z^{\star},W_{t}^{\star}]$ from each network is well-established, $Y_{t}^{\star}$ is only identifiable up to an unknown orthogonal transformation under a single-network model. As a result, $Z^{\star}$ cannot be directly obtained. To address this issue, we leverage joint information of $Y_{t}^{\star}$’s and develop a novel and efficient spectral method below. To motivate our proposed construction, we first introduce an algebraic result in an oracle scenario below.
@@
    there exist $1\leqslant t<s\leqslant T$ such that columns of $Z^{\star},W_{t}^{\star},W_{s}^{\star}]$ are linearly independent.
Let $\mathcal{T}=\{(t,s):1\leqslant t<s\leqslant T\text{ and Assumption (ii) holds for }(t,s)\}$. For $(t,s)\in\mathcal{T}$,
@@
Assumption (i) in Proposition 2 posits an oracle situation where we can accurately estimate each $Y_{t}^{\star}$ individually up to the orthogonal group transformation $\mathcal{O}(d_{t})$. Assumption (ii) in Proposition 2 is the same as Assumption (ii) in Proposition 1. Eq. (3) provides the theoretical foundation for identifying $Z^{\star}$, which is unachievable when only investigating each $Y_{t}$ individually. However, (3) may not be directly applicable to estimating $Z^{\star}$ in practice because an estimator $\mathring{Y}_{t}$ obtained in Step (a) may not exactly satisfy $\mathring{Y}_{t}\mathring{Y}_{t}^{\top}=Y_{t}^{\star}Y_{t}^{\star\top}$, and $\mathcal{T}$ relies on unknown true parameters.
To address these issues, we develop practical procedures based on spectral information of the estimates $\mathring{Y}_{t}$’s in Step (a). We first estimate $\mathcal{T}$. Let $\sigma_{i}(B)$ denote the $i$-th largest singular value of a matrix $B$, and define $\mathcal{R}_{i,j}(B)=\sigma_{i}(B)/\sigma_{j}(B)$. Then we have
@@ 3.3 Likelihood-based refinement
$\ell(Z,W)=\sum_{t=1}^{T}\sum_{1\leqslant i\leqslant j\leqslant n}l\big{(}\langle z_{i},z_{j}\rangle+\langle w_{t,i},w_{t,j}\rangle;A_{t,ij}\big{)},$ (5)
where $l(\theta;x)=\log p(\,x\mid\theta\,)$ and $W=[W_{1},\ldots,W_{T}]$. Our refinement procedure first conducts the projected gradient descent (Chen and Wainwright,, 2015), primarily using first-order derivatives of the likelihood function that can be efficiently computed. Specifically, the algorithm descends parameters along their gradient directions with pre-specified step sizes and then projects updated estimates to pre-specified constraint sets for parameters. The detailed steps are summarized in Algorithm 2 with notations explained in Section 3.4. Let $\check{Z}$ and $\check{W}_{t}$’s denote the estimates after sufficient iterations of the projected gradient descent.
Then the refinement procedure further constructs a second-order update, which utilizes efficient influence function and provides convenience for theoretically establishing desired statistical error rate. More discussions on its theoretical usefulness are provided in Remark D.6 of the Supplementary Material. For the ease of introducing formula below, we denote a vectorization of all the entrywise parameters of $Z,W_{1},\ldots,W_{T}]$ as
@@
where $k_{\operatorname{sum}}=k_{1}+\cdots+k_{T}$, and similarly define $\check{v}$ as the vectorization of $\check{Z},\check{W}_{1},\ldots,\check{W}_{T}]$. Then $\ell(Z,W)=\ell(v)$, and let $\dot{\ell}(v)$ denote the partial derivative of $\ell(v)$ with respect to $v$. We construct a second-order update as
@@
Our proof shows $I(\check{v})$ in (7) is singular, necessitating the use of pseudo inverse in (7). Intuitively, this singularity arises from the unidentifiability of latent factors $Z$ and $W_{t}$ (Little et al.,, 2010). It makes the theoretical analysis challenging and conclusions for classical one-step estimator (Van der Vaart,, 2000) or a single network Xie and Xu, (2023) inapplicable; more details are discussed in Remark D.5 of the Supplementary Material. Moreover, (7) is simultaneously constructed for $1+T$ groups of latent factors $Z,W_{1},\ldots,W_{T}]$. This markedly differs from He et al., (2025) that constructs one-step estimator for shared latent factors only and requires network-specific parameters to be identifiable. As a result, our theoretical analysis pioneers a new approach that jointly analyzes all the parameters in (6) and tackles inherent singularity of the information matrix from multiple groups of latent factors. More details are given in Section D.5.4 of the Supplementary Material.
@@ 3.4 Algorithm details
10Construct $\hat{v}$ by (7) with $(\check{Z},\check{W})=(Z^{R},W^{R})$, and let $\hat{Z},\hat{W}$] be its matrix version.
@@
In Algorithm 1: The threshold $\tau_{1}$ is chosen to estimate $\mathcal{T}$. As Section 3.2 suggests, the choice of $\tau_{1}$ depends on statistical properties of the input initial estimates $\mathring{Y}_{t}$. In this paper, we construct $\mathring{Y}_{t}$ by the method in Section 3.1, whose estimation error is established in Section A.2 of the Supplementary Material. To achieve theoretical guarantee, it suffices to set the threshold to satisfy $1\ll\tau_{1}\lesssim\sqrt{\log n}$. Our numerical implementation sets $\tau_{1}=\sqrt{2\log n}$.
In Algorithm 2: During the gradient descent, we let $\dot{\ell}_{Z}(Z,W)\in\mathbb{R}^{n\times k}$ and $\dot{\ell}_{W_{t}}(Z,W)\in\mathbb{R}^{n\times k_{t}}$ represent the partial derivatives of $\ell(Z,W)$ with respect to matrices $Z$ and $W_{t}$, respectively. Let $\mathcal{P}_{\mathcal{C}}(\cdot)$ denote a projection operator given a pre-specified constraint set $\mathcal{C}$, and we consider two constraint sets
@@
for $Z$ and $W_{t}$, respectively. Eq. (8) constraints the two-to-infinity norms of $Z$ and $W_{t}$ to be bounded by $M_{1}$, which also corresponds to the constraints on the true parameters $(Z^{\star},W_{t}^{\star})$ in Condition 2 below. The projection steps and second-order update in Algorithm 2 are required for the convenience of the proof, and skipping them can yield equally good numerical performance. More details are discussed in Section G.2 of the Supplementary Material. In our numerical implementation, we use Barzilai-Borwein step sizes (Barzilai and Borwein,, 1988) and set $R=1000$. To establish theoretical results in Section 4, Condition 1 below summarizes the requirements on the hyperparameters in Algorithms 1 and 2.
@@
Assume the hyperparameters in Algorithms 1 and 2 satisfy: (i) $1\ll\tau_{1}\lesssim\sqrt{\log n}$. (ii) Step sizes $\eta_{Z}=\eta/(nT)$ and $\eta_{W}=\eta/n$ for a sufficiently small constant $\eta>0$. (iii) Constraint sets $\mathcal{C}_{Z}$ and $\mathcal{C}_{W_{t}}$ are chosen as in (8). (iv) Number of iterations $R\gg\log(nT)$.
@@ 3.5 Related works
The proposed estimation scheme in Section 3 shares connections with, yet notably differs from, existing methods. The shared space hunting Algorithm 1 is developed based on spectral properties of latent factors. While spectral-based analysis is common in network studies (Rohe et al.,, 2011; Chatterjee,, 2015; Arroyo et al.,, 2021), our procedure utilizes unique properties of multiple networks to separate the shared and distinct latent spaces, and thus cannot be directly implied by existing methods. The likelihood refinement Algorithm 2 utilizes the projected gradient descent, which is known to be computationally efficient for maximizing the non-convex likelihood functions for both single network (Ma et al.,, 2020) and multiple networks (Zhang et al.,, 2020). Nevertheless, the oracle estimation error rates under our model cannot be directly attained through existing technical developments of those methods. To address the theoretical challenges, Algorithm 2 further constructs a novel second-order update estimator (7) for multiple heterogeneous networks. The construction is motivated from the classical one-step estimator, typically obtained through a linear approximation of score equation (Van der Vaart,, 2000). Similar idea has also been recently used to construct efficient estimators for latent factors in network analysis, including the random dot product model for a single network (Xie and Xu,, 2023) and a semiparametric longitudinal network model (He et al.,, 2025). However, our model (1) substantially differs from those models in that it embraces a general edge distribution and also allows distinct latent factors over multiple networks. The extended flexibility posits unique challenges in identifying shared spaces and obtaining efficient estimation errors, necessitating novel developments of efficient score updating formula and theory.
The model (1) is motivated from MacDonald et al., (2022), which similarly considers (1) and assumes that part of the latent vectors are shared across the networks, i.e., $y_{t,i}=(z_{i}^{\top},w_{t,i}^{\top})^{\top}$. The difference is that in their work $\Theta_{t,ij}=\kappa(y_{t,i},y_{t,j})$ with $\kappa(x,y)=x^{\top}\mathrm{I}_{p,q}y$, where $\mathrm{I}_{p,q}$ represents the block-diagonal matrix with $\mathrm{I}_{p}$ and $\mathrm{I}_{q}$ on the upper left and lower right blocks, respectively. Although a more general similarity function is proposed, their theoretical analysis is restricted to Gaussian distribution with identity link, which can be inappropriate for network data with binary edges or count edges. Alternatively, we will next establish theoretical results that encompass various exponential family distributions. Please also see Remark 4 for more discussions on the comparison of technical conditions and results.
@@
Model Link $\mu(\theta)$ ER-Shared ER-Individual Constraints on $(n,T)$
Arroyo et al., (2021) $\theta$ $O_{p}\biggr{(}\frac{1}{n}+\frac{1}{T}\biggr{)}$ - -
MacDonald et al., (2022) $\theta$ $O_{p}\biggr{(}\frac{1}{T}\biggr{)}$ $O_{p}(1)$ $T=o(n^{{1}/{2}})$
Zhang et al., (2020) $\operatorname{expit}(\theta)$ $O_{p}\biggr{(}1+\frac{1}{T}\biggr{)}$ - $T=O(n)$
He et al., (2025) $\exp(\theta)$ $O_{p}\biggr{(}\frac{1}{n}+\frac{1}{T}\biggr{)}$ $O_{p}(1)$ $\mathrm{polylog}(T)=O(n)$
This work General $O_{p}\biggr{(}\frac{1}{n}+\frac{1}{T}\biggr{)}$ $O_{p}(1)$ $\mathrm{polylog}(T)=O(n)$
(Error rates above are simplified to emphasize their orders with respect to $n$ and $T$ up to logarithmic factors and also rescaled to be comparable. “-” represents inapplicable information. See more details in Section F of the Supplementary Material.)
Besides the studied model (1), there are various other models extracting node-wise latent embeddings from multiple networks. One research line extends the random dot product graphs (Young and Scheinerman,, 2007; Athreya et al.,, 2018) to multilayer networks (Arroyo et al.,, 2021; Zheng and Tang,, 2022). These models typically corresponds to an identity link function in (2), and thus developments cannot directly address the challenges under models with general non-linear link functions. Another class of models adopt non-linear canonical link functions (Hoff et al.,, 2002; Ma et al.,, 2020) and develop extensions for multiple networks (Zhang et al.,, 2020; MacDonald et al.,, 2022; He et al.,, 2025). Our developments under (1) advance the literature by establishing a unified theoretical framework that can encompass a variety of general link functions and allow heterogeneous latent factors with complex relationships. We summarize properties and results under the related models in Table 1, where our theory will be formally established in Section 4. Notably, oracle error rate of the shared parameters in this paper is consistent with earlier work in Arroyo et al., (2021), MacDonald et al., (2022), and He et al., (2025).
Targeting at distinct properties and challenges, researchers have also proposed other models for multiple networks. Examples include multilayer stochastic block models for community detection (Han et al.,, 2015; Paul and Chen,, 2016; Lei and Lin,, 2023), and tensor-decomposition-based models for effective dimension reduction (Lyu et al.,, 2023; Zhang and Wang,, 2023), etc. Although these models may also incorporate shared and heterogeneous structures across multiple networks, their analytical objectives and challenges differ from our focus on estimating continuous node-wise latent embeddings and thus are not directly compared in Table 1; see more discussions in Section F of the Supplementary Material.